Nash Inc. issues 500 shares of $10 par value common stock and 100 shares of $100 par value preferred stock for a lump sum of $110,000.

(a) Prepare the journal entry for the issuance when the market price of the common shares is $180 each and market price of the preferred is $225 each.
(b) Prepare the journal entry for the issuance when only the market price of the common stock is known and it is $190 per share.


(Round answers to 0 decimal places, e.g. $1,225. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

No.

Account Titles and Explanation

Debit

Credit

(a)

enter an account title for case A

enter a debit amount

enter a credit amount

enter an account title for case A

enter a debit amount

enter a credit amount

enter an account title for case A

enter a debit amount

enter a credit amount

enter an account title for case A

enter a debit amount

enter a credit amount

enter an account title for case A

enter a debit amount

enter a credit amount

(b)

enter an account title for case B

enter a debit amount

enter a credit amount

enter an account title for case B

enter a debit amount

enter a credit amount

enter an account title for case B

enter a debit amount

enter a credit amount

enter an account title for case B

enter a debit amount

enter a credit amount

enter an account title for case B

enter a debit amount

enter a credit amount
Answer #1
Dear Student
Thank you for using homeworklib
Please give thumbs up if you are satisfied
a)
Fair value of common = 500*180          90,000.00
Fair value of Preferred = 100*225          22,500.00
       112,500.00
Allocated to common = 90,000/112,500*110,000          88,000.00
Allocated to Preferred = 22,500/112,500*110,000          22,000.00
Total Allocation        110,000.00
Debit Credit
Cash Dr        110,000.00
To Common Stock=500*10            5,000.00
To Paid-in Capital in Excess of Par - Common Stock = 88000-5000          83,000.00
To Preferred Stock = 100*100          10,000.00
To Paid-in Capital in Excess of Par - Preferred Stock = 22000-10000          12,000.00
b)
Lumpsum Receipt        110,000.00
Allocated to Common = 500*190          95,000.00
Balance allocated to preferred          15,000.00
Debit Credit
Cash Dr        110,000.00
To Common Stock=500*10            5,000.00
To Paid-in Capital in Excess of Par - Common Stock = 95000-5000          90,000.00
To Preferred Stock = 100*100          10,000.00
To Paid-in Capital in Excess of Par - Preferred Stock = 15000-10000            5,000.00